OpenStudy (anonymous):

Twenty pounds of a salt-water solution contains fifteen percent salt. How much water must be added to weaken the solution to twelve percent salt?

OpenStudy (amistre64):

20(.15) + w(0) = (20+w)(.12)

OpenStudy (amistre64):

3 = 2.4 +.12w 3 - 2.4 = .12w .6 --- = w .12 w = 60/12 = 5

OpenStudy (anonymous):

thaksnada

OpenStudy (amistre64):

yep :)